# General Chemistry Exam with Artificial Intelligence Jan Cvengros,a and Gerd Kortemeyerb†

### v1

@@ 6 Conclusions
In this study, we explored the effectiveness of AI-assisted grading for open-ended, handwritten general chemistry exams, aiming to balance efficiency with grading quality in large-enrollment courses. By comparing AI-generated scores to expert human graders, we demonstrated that carefully calibrated AI systems can substantially reduce grading workloads without sacrificing accuracy. Our analysis indicated that AI reliably identifies fully correct answers, significantly streamlining grading when used selectively.
@@ Acknowledgements
We would like to thank the students who chose to participate in our study. We would also like to thank Anna Kortemeyer for her help with data entry.

### v2

@@ 7 Conclusions
In this study, we explored the effectiveness of AI-assisted grading for open-ended, handwritten general chemistry exams, aiming to balance efficiency with grading quality in large-enrollment courses. By comparing AI-generated scores to expert human graders, we demonstrated that carefully calibrated AI systems can substantially reduce grading workloads without sacrificing accuracy. Grading accuracy could be further improved by averaging multiple, independent runs. Our analysis indicated that deploying confidence filtering yields more dependable identification of fully correct answers, significantly streamlining grading when used selectively.
@@
Ethical Considerations
This study was approved by the ETH Zurich Ethics Commission (IRB) as protocol EK 2023-N-286. Students gave voluntary informed consent for participation.
@@ Acknowledgements
We would like to thank the students who chose to participate in our study. We would also like to thank Anna Kortemeyer for her help with data entry. We would also like to thank the reviewers of this journal for their insightful and constructive criticism.